OpenStudy (anonymous):

How many solutions does the equation −2|x − 3| + 3 = 4 have? I think it is one because, I got, x = 2.5 and x = 3.5 −2|x − 3| + 3 = 4 I plugged them in and only the 2.5 equals 4.

jimthompson5910 (jim_thompson5910):

−2|x − 3| + 3 = 4 −2|x − 3| = 4-3 −2|x − 3| = 1 |x − 3| = 1/(-2) |x − 3| = -1/2 Can you see any issues here?

OpenStudy (anonymous):

you need the first part of the equation to equal 1 for it to work, but it's a negative number multiplied by a positive one, which will always yield a negative number, therefore, there are no solutions to this equation